OpenStudy (anonymous):

Find the Orthogonal Trajectories of the the curve... y^2=kx^3

OpenStudy (anonymous):

first differentiate the expretion: 2yy'=3kx^2 now substitute y'=-1/y': -2y/y'=3kx^2 separate variables and integrate: dy/2y=-(1/3k) x^-2 1/2Ln|y|=(1/3k)x^-1 + Ln C y^(1/2)=Ce^1/(3kx) y=Ce^2/(3kx)

OpenStudy (anonymous):

Thanks! :D